randy19234:

kayla used her credit card to purchase a new television for $487.89. she can pay off up to $175 per month. the card has an annual rate of 23.5% compunded monthly. how much will she pay in interest?

Angle:

@BRACEFACEJAH you are incorrect, the annual interest rate is 23.5% which means that you should be paying 1.958333% in interest per month

Angle:

Otherwise, the process is correct. Month 1: Owed Amount = 487.89 Kayla Pays = 175 Amount Left = 312.89 1.958333% of 312.89 = 6.127 this gets added to the owed amount Month 2: Owed Amount = 312.89 + 6.13 = 319.02 Kayla Pays = 175 Amount Left = 144.02 1.958333% of 144.02 = 2.82 this gets added to the owed amount Month 3: Owed Amount = 144.02 + 2.82 = 146.84 Kayla Pays = 175 Amount Left = 0 ---- Amount paid in interest = 6.13 + 2.82